# Binaural Target Speaker Extraction using HRTFs and a Complex-Valued Neural Network

### v1

@@ IV-E Results
In Fig. 3, we show histograms of the SI-SDR, PESQ, and STOI metrics before and after applying our TSE algorithm in the anechoic scenario. All three distributions shift markedly to the right, demonstrating substantial gains across signal quality, perceptual quality, and intelligibility. This confirms the algorithm’s effectiveness in isolating the desired speaker.
In Table I, we compare our method against competing algorithms across all datasets under both anechoic and reverberant conditions. Our approach achieves significant gains on every metric except PESQ, where the BLCMV beamformer slightly outperforms the proposed method in anechoic case. This may be attributed to its distortionless response design criterion.
When comparing the two additional variants of our model, denoted as Bi-TSE-DOA and Bi-TSE-HRTF-RI in Table I, we observe that, in contrast to the HRTF-guided model, the DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and is unable to preserve the binaural cues. We evaluated this model under reverberant conditions to emphasize that the HRTF is intrinsically tied to our problem formulation. The superior performance of the HRTF-guided network supports these findings. On the other end, the RI model achieves comparable SI-SDR and binaural cue preservation to the complex-valued model but suffers from lower output quality, as shown by its lower PESQ score in the anechoic setting.
@@
Remarkably, our approach demonstrates cross-dataset generalizability: although our model was trained exclusively on WSJ0-CSR1, it performs exceptionally well on other corpora, including a French dataset. We attribute this generalization to the absence of speaker-specific priors in our approach.
Figure 4: A polar plot showing the SI-SDR of the output signals for each DOA, with the desired speaker located at $\theta_{d}=-75^{\circ}$ (red) and the interfering speaker at $\theta_{i}=-10^{\circ}$ (blue). The bold circle represents the 0 dB SI-SDR threshold.
When inspecting the extracted signals, we found that the model effectively enhances the target speaker while nearly eliminating the interferer. To quantify its spatial resolution, we devised a “scanning” procedure: we iteratively extract signals using HRTFs corresponding to azimuth angles in $[-90^{\circ},90^{\circ}]$ at a fixed elevation angle and compute the SI-SDR of each extraction against the true target. Plotting these values yields a beam-pattern-like diagram (Fig. 4), which shows a pronounced peak at the correct DOA. Additionally, the width of the main lobe indicates that the model remains effective even when the HRTF clue deviates slightly from the true HRTF. The proposed method’s ability to preserve binaural cues is also evaluated and compared with the BLCMV, where applicable. Table I presents the mean ILD and ITD deviations. Figure 5 depicts the p.d.f.s of both cues for the target and extracted signals, showing a close match and confirming faithful spatial cue preservation.
(a) Mixture
(b) $\hat{\tilde{\bm{s}}}_{d}$
(c) $\hat{\tilde{\bm{s}}}_{i}$
Figure 5: The p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=90^{\circ}$, $\theta_{i}=15^{\circ}$, and $\phi_{d,i}=0^{\circ}$. Graphs produced by [31].
@@ V Conclusions
This paper presents a speaker‑independent TSE method that leverages the listener’s HRTFs in both anechoic and reverberant settings, resulting in a personalized solution for the proposed problem. Through cross‑dataset evaluations, we demonstrate its robust generalization across diverse corpora. Our approach not only delivers high‑quality speech but also faithfully preserves the binaural character of the signal. Comparisons with other binaural TSE techniques underscore the importance of spatial cues tied to the listener’s HRTF for accurate and consistent target‑speaker extraction under varying acoustic conditions.

### v2

@@ IV-E Results
In Table I, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions. In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE. Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
@@
Remarkably, both DNN-based methods demonstrate strong cross-dataset generalizability. Although the models were trained exclusively on WSJ0-CSR1, they perform well on other corpora, including a French dataset. We attribute this generalization to the absence of speaker-specific priors and to the reliance on a known HRTF database.
[…]
BLCMV Beamformer 1.7 >90>90 >150>150
BDE-BiTSE 1.0667 30 90
Bi-TSE-HRTF 0.0 -60 0
Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI. The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues, especially under reverberant conditions, underscoring the central role of HRTFs in our formulation. By contrast, the RI model achieves SI-SDR and binaural cue preservation comparable to the complex-valued model in the anechoic case, but suffers from reduced perceptual quality (lower PESQ) and fails to maintain cue preservation under reverberation.
Figure 3 shows the joint ITDILD p.d.f.s across methods. The close alignment of our model’s distribution with that of the target underscores its effectiveness in maintaining spatial cues. In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations. The dereverberation effect is also apparent, as the joint ITD-ILD p.d.f. becomes more concentrated, indicating dominance of the direct path in the output of the proposed method. Overall, it is evident that the proposed approach achieves this more effectively than competing techniques. Table LABEL:tab:delta_thata further highlights the spatial consistency observed in Fig. 3. Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec. IV-C. The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
Fig. 4: A polar plot of SI-SDR versus DOA, with the target at $\theta_{d}=-75^{\circ}$ (red) and the interferer at $\theta_{i}=-10^{\circ}$ (dashed blue). Stars mark true DOAs; the dotted arc denotes the 0,dB threshold.
Finally, to quantify the spatial resolution of our model, we employ a “scanning” procedure: signals are iteratively extracted using HRTFs at azimuths within $90^{\circ},90^{\circ}]$ (with fixed elevation), and the SI-SDR of each extraction is computed against the target. Plotting these values produces a beam-pattern-like diagram (Fig. 4) that exhibits a clear peak at the true DOA. The width of the main lobe further characterizes the model’s robustness when the HRTF cue deviates slightly from the ground-truth.
@@ V Conclusions
This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, yielding a personalized solution to the task at hand. Through cross-dataset evaluations, we demonstrate robust generalization across diverse corpora. The proposed approach not only produces high-quality speech but also faithfully preserves the binaural characteristics of the signal. Comparative analyses with existing binaural TSE methods reveal substantial improvements in spatial consistency, while preserving high speech quality. These findings highlight the critical role of the listener’s HRTF in anchoring spatial cues, enabling accurate and consistent target-speaker extraction under a wide range of acoustic conditions.